OpenStudy (ttp):

4.A quality control inspector is in charge of inspecting drill bits that are produced on an assembly line. The diameter of a drill bit can be modeled by the equation |x − 0.525| = 0.01. Show your work in solving the equation to determine the minimum and maximum diameters of the drill bits.

OpenStudy (anonymous):

whoa should be \[|x-0.525|\leq 0.01\]

OpenStudy (ttp):

It's not.

OpenStudy (anonymous):

well it if is not, it should be let me guess, another FLVS question?

OpenStudy (anonymous):

in any case it means \(x\) is within \(0.01\) units of \(0.525\) solve \[-0.01\leq x-0.525\leq 0.01\] by adding \(0.525\) all the way across

OpenStudy (ttp):

I have any answer could you check it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you should have two numbers, right?

OpenStudy (ttp):

x - 0.525| = 0.01 x - 0.525 = 0.01 x = 0.01 + 0.525 x = 0.535 -| x - 0.525| = 0.01 -x + 0.525 = 0.01 -x = 0.01 - 0.525 -x = 0.515 x = - 0.515

OpenStudy (anonymous):

top one is right, bottom not

OpenStudy (anonymous):

it is not possible for \(-|x-0.525|\) to be equal to \(0.01\) because the absolute values is positive think you got confused on the second one but we can fix is quickly

OpenStudy (anonymous):

when you solve \[|x-0.525=0.01|\] you do have to write two equations, but those equation so not have absolute values in them the two equations you solve are \[x-0.525=0.01\] which was your first one, and also \[x-0.525=-0.01\]

OpenStudy (anonymous):

typo there, i meant when you solve \[|x-0.525|=0.01\]

OpenStudy (ttp):

Could you help me get the right answer on the second?

OpenStudy (anonymous):

the second is \[x-0.525=-0.01\] solve just like the first, by adding \(0.525\) to both sides, gets \[x=0.525-0.01\]

OpenStudy (anonymous):

so two answers are \[0.535\] and \[0.515\] i.e. \(0.01\) units above and below \(0.525\)

OpenStudy (ttp):

I had that?

OpenStudy (anonymous):

hmm no you had it negative -x + 0.525 = 0.01 -x = 0.01 - 0.525 -x = 0.515 x = - 0.515

OpenStudy (anonymous):

actually the first line there \[-x+0.525=.0.1\\ -x=0.01-0.525\] are right

OpenStudy (ttp):

x - 0.525| = 0.01 x - 0.525 = 0.01 x = 0.01 + 0.525 x = 0.535 -x + 0.525 = 0.01 -x = 0.01 - 0.525 -x = 0.515 x = - 0.515 so these are correct?

OpenStudy (anonymous):

yes

OpenStudy (ttp):

?

OpenStudy (ttp):

Hello?

OpenStudy (anonymous):

yes